# Neuronal Spike Trains as Functional-Analytic Distributions: Representation, Analysis, and Significance

### v1

@@ 5 Integral notation for the Dirac delta
While seemingly unnecessary, this notation has a number of important consequential advantages. It allows us to deal with smooth functions and spikes in a single functional framework, and maintains consistency when working with convolutions and integration by parts. We show this below.
@@ 6 How the Dirac Delta Arises as a Limit of Ordinary Functions
Notice how the area reamins equal to 1, since
@@
(A necessary technical note: Following standard functional analysis notation, we write $\mathcal{D}^{\prime}$ to indicate the space of (Schwartz) distributinos. The space of distributions $\mathcal{D^{\prime}}(\mathbb{R})$ is the dual space to the space of test functions: $\mathcal{D}(\mathbb{R})=C_{c}^{\infty}(\mathbb{R})$. $C_{c}^{\infty}(\mathbb{R})$ is formally a vector space, and the dual space of distributions is the space of linear functionals. So writing $\rho_{\varepsilon}(t-a)\xrightarrow{\ \mathcal{D}^{\prime}\ }\delta_{a}$ in Equation 6.1 implies that in the limit the functions converge in the sense of distributions.)
@@ 7 Putting it all Together: The Dynamics of Spike Trains
In this last section, we apply the framework developed above to the dynamics of spike trains. Fundmentally, the purpose of this formalism is not just to represent spikes more accurately, but to enable mathematically rigorous and consistent modeling and analysis of neuronal dynamics that cannot be treated coherently without it.
@@ 7.1 Linear summation of spike distributions
Whereas, Equation 4.2 satisfies the operational requirment we set up in Equation 2.1 for the single value case, Equation 7.3 equivalently does so for a linear combination of functionals.
@@ 7.2 Local finiteness and well-defined spike trains
We address a technical consideration to ensure that the spike train representation is mathematically well-defined. Because distributions are defined by their action on compactly supported test functions, the only way an infinite sum of spike events could fail to be well-defined is if infinitely many spikes contribute within a finite time window. This condition is avoided by assuming local finiteness, that is, only finitely many spikes may occur in any bounded interval of time, even though the spike train may contain infinitely many events overall, athematically speaking at least. In practice, of course, any spike train is necessarily finite. But this requirment ensures mathematical consistency.
Intentionally, test functions $\varphi\in C_{c}^{\infty}(\mathbb{R})$ have compact support (c.f. Section s 4.3 and 4.4). This ensures that there exists a finite interval $[a,b]$ such that $\varphi(t)=0$ outside $[a,b]$. Because $t\in\mathbb{R}$, the support of $\varphi$ takes the infinite sum
@@
since every term $t_{i}^{k}\notin[a,b]$ is zero. The only remaining condition we need to impose is that that there not be infinitely many spike times inside $[a,b]$. So we assume local fititeness such that
@@
This guarantees that for every test function only finite many nonzero terms (spikes) appear in the sum, making the sum finite and precise. Otherwise, the sum becomes a conditionally convergent (or divergent) infinite series due to Riemann rearrangement, whereby any conditionally convergent infinite series of real numbers can be rearranged to converge to any value, or to diverge. From a neurophysiological and neural dynamics perspective, this constraint is, of course, reasonable, since we would never record an infinite number of spikes in a finite time window.
@@ 7.3 Convolution and spike train dynamics
We want to construct a spike train $s_{i}$ such that each spike at time $t_{i}^{k}$ produces a stereotyped postsynaptic response waveform kernel $K$ that gets triggered at that time. Consistent with the formalism we laid out for Equation 2.1, the subscript $i$ indexes individual neurons, while the superscript $k$ indexes individual spike events within the spike train of neuron $i$. Leveraging the linear summation of multiple occuring spiking events, we can write an expression for synaptic current $I_{syn}$ as
@@
Neurophysiologically, $K$ here represents the elemetary postsynaptic current contributed by each presynaptic neuron on the postsynaptic neuron $i$. $I_{syn}(t)$ is the linear supperposition of individual currents, i.e., the total synaptic current. For example, a standard form for a postsynaptic current kernel for ionotropic $Cl^{-}$ gamma-aminobutyric acid (GABA) and $Na^{+}$ $\alpha$-amino-3-hydroxy-5-methyl-4-isoxazolepropionic acid (AMPA) ion channels is
$K(t)=g_{\max}\,\bigl(e^{-t/\tau_{d}}-e^{-t/\tau_{r}}\bigr)H(t)$
@@
Extending this to a series of spikes, i.e., a spike train $s_{i}(\tau)=\sum_{k}\delta(\tau-t_{i}^{k})$, assuming linearitiy (Sections 4.2 and 7.1) and local finiteness (Section 7.2), we compute the convolution in the same way (dropping the integration limits for notational simplicity):
@@
In the context of convolution, the Dirac delta acts as a form of time translation. When a synaptic kernel $K$ is convolved with a time-shifted delta $\delta(t-a)$, the delta essentially translates the entire shape of the kernel along the timeline to exactly $t=a$. This functionally transforms a complex integral into a simple copy-paste operation: for every spike time in the collection of events, the distributiional operation pastes a copy of the synaptic response into the total current, allowing discrete events to be encoded into continuous-time dynamics.
@@ 7.4 Distributional differentiation of spike trains
We end the paper by showing another example that emphasizes the unique gains of a distribution theory formalism to neuronal spike trains. Although a spike train has no pointwise derivative, its derivative is well-defined in the distributional sense. By definition, the derivative $s_{i}^{\prime}$ acts on a test function $\varphi$ as $\langle s_{i}^{\prime},\varphi\rangle=-\langle s_{i},\varphi^{\prime}\rangle$, producing $\sum_{k}\varphi^{\prime}(t_{i}^{k})$. In other words, differentiation of a spike train does not measure changes in the spikes themselves, which has no meaning, but rather local changes in the test function at the event times. We will develop this intuition carefully.
@@
A spike train is a set of instantaenous events, so what does it even mean to differentiate a set like that? When we differentiate ordinary functions, derivatives have phyical and geometric interpretations, such as the rate of change, or slope. But distributions like the spike train are not ordinary functions. They have no meaning or values between events, or concepts like slope. This necessitates a change in perspective that ultimately results in a powerful set of mathematical ideas and tools to analyze and understand (in our case) very real world physical and physiological properties of neural signaling and dynamics.
@@
This shift in perspective involves asking not how do we differentitae the distribution or what that might mean. The spike train cannot be directly differentiated. Instead, we differentiate the test functions and ask how does the spike train respond.
From the operational requirement set in Equation 2.1,
@@
with $s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k})$ (Equation 4.4). The spike train samples the test function at the event times.
@@
This definition is consitent with the ordinary notion of differentiation if $T$ were a smooth function, and preserves integration by parts, but avoids pointwise evaluation. We look at this more deeply (and why the negative sign in the definition needs to be there) below.
@@
The intepretation of Equation 7.10 is that the original spike train $s_{i}$ samples values of the test function $\varphi(t)$ at the event times, while the derivative of the spike train, $s_{i}^{\prime}$, samples values of the derivative of the test function $\varphi^{\prime}(t)$. In other words, $s_{i}^{\prime}$ samples values of the slope of $\varphi(t)$. This means that the derivatives of the spike train (by definitiion) measures how the test function changes at the event times.
[…]
Two closing technical comments: First, if $\varphi(t)$ is not differentiable, then $\varphi^{\prime}(t)$ would be undefined, making $s_{i}^{\prime}$ undefined (directly due to its definition). So this operation necessitates all the work we did above to justify why $\varphi(t)\in C_{c}^{\infty}$, and as a result needs to be smooth (in a mathematically formal sense).
Second, the minus sign in Equation 7.9 results directly from integration by parts. It is worth cosidering why it arrises in some detail, because it provides a deep intuative insight.
@@
Notice how ther right hand side of Equation 7.13d is (by definition) $\langle T^{\prime}_{f},\varphi\rangle$ (c.f. Equation 7.11). The left hand side is Equation 7.12: $\langle T_{f},\varphi^{\prime}\rangle$. Therefore,

### v2

@@ 5 Integral notation for the Dirac delta
While seemingly unnecessary, this notation has a number of important consequential advantages. It allows us to deal with smooth functions and spikes in a single notational framework, and maintains consistency when working with convolutions and integration by parts. We show this below.
@@ 6 How the Dirac Delta Arises as a Limit of Ordinary Functions
Notice how the area remains equal to 1, since
@@
Following standard functional analysis notation, we again write $\mathcal{D}^{\prime}$ to indicate the space of (Schwartz) distributions. The space of distributions $\mathcal{D^{\prime}}(\mathbb{R})$ is the dual space to the space of test functions: $\mathcal{D}(\mathbb{R})=C_{c}^{\infty}(\mathbb{R})$. $C_{c}^{\infty}(\mathbb{R})$ is formally a vector space, and the dual space of distributions is the space of linear functionals. So writing $\rho_{\varepsilon}(t-a)\xrightarrow{\ \mathcal{D}^{\prime}\ }\delta_{a}$ in Equation 6.1 implies that in the limit the functions converge in the sense of distributions.
@@ 7 Putting it all Together: The Dynamics of Spike Trains
In this section, we apply the framework developed above to the dynamics of spike trains. Fundamentally, the purpose of this formalism is not just to represent spikes more accurately, but to enable mathematically rigorous and consistent modeling and analysis of neuronal dynamics that cannot be treated coherently without it. In Section 7 here, we derive some theoretical results. In Section 8 below, we apply these results to the analysis of a simple reciprical two neuron circuit.
@@ 7.1 Linear summation of spike distributions
Whereas, Equation 4.2 satisfies the operational requirement we set up in Equation 2.1 for the single value case, Equation 7.3 equivalently does so for a linear combination of functionals.
@@ 7.2 Local finiteness and well-defined spike trains
We next address a technical consideration to ensure that the spike train representation is mathematically well-defined. Because distributions are defined by their action on compactly supported test functions, the only way an infinite sum of spike events could fail to be well-defined is if infinitely many spikes contribute within a finite time window. This condition is avoided by assuming local finiteness, that is, only finitely many spikes may occur in any bounded interval of time, even though the spike train may contain infinitely many events overall, mathematically speaking at least. In practice, of course, any spike train is necessarily finite. But this requirement ensures mathematical consistency.
Intentionally, test functions $\varphi\in C_{c}^{\infty}(\mathbb{R})$ have compact support (c.f. Sections 4.3 and 4.4). This ensures that there exists a finite interval $[a,b]$ such that $\varphi(t)=0$ outside $[a,b]$. Because $t\in\mathbb{R}$, the support of $\varphi$ takes the infinite sum
@@
since every term $t_{i}^{k}\notin[a,b]$ is zero. The only remaining condition we need to impose is that there not be infinitely many spike times inside $[a,b]$. So we assume local finiteness such that
@@
This guarantees that for every test function only finitely many nonzero terms (spikes) appear in the sum, making the sum finite and precise. Otherwise, the sum becomes a conditionally convergent (or divergent) infinite series due to Riemann rearrangement, whereby any conditionally convergent infinite series of real numbers can be rearranged to converge to any value, or to diverge. From a neurophysiological and neural dynamics perspective, this constraint is, of course, reasonable, since we would never record an infinite number of spikes in a finite time window.
@@ 7.3 Convolution and spike train dynamics
We want to construct a spike train $s_{i}$ such that each spike at time $t_{i}^{k}$ produces a stereotyped postsynaptic response waveform kernel $K$ that gets triggered at that time. Consistent with the formalism we laid out for Equation 2.1, the subscript $i$ indexes individual neurons, while the superscript $k$ indexes individual spike events within the spike train of neuron $i$. Leveraging the linear summation of multiple occurring spiking events, we can write an expression for synaptic current $I_{syn}$ as
@@
Neurophysiologically, $K$ here represents the elementary postsynaptic current contributed by each presynaptic neuron on the postsynaptic neuron $i$. $I_{syn}(t)$ is the linear superposition of individual currents, i.e., the total synaptic current. For example, a standard form for a postsynaptic current kernel for ionotropic $Cl^{-}$ gamma-aminobutyric acid (GABA) and $Na^{+}$ $\alpha$-amino-3-hydroxy-5-methyl-4-isoxazolepropionic acid (AMPA) ion channels is
$K(t)=g_{\max}\,\bigl(e^{-t/\tau_{d}}-e^{-t/\tau_{r}}\bigr)H(t)$ (7.5)
@@
Extending this to a series of spikes, i.e., a spike train $s_{i}(\tau)=\sum_{k}\delta(\tau-t_{i}^{k})$, assuming linearity (Sections 4.2 and 7.1) and local finiteness (Section 7.2), we compute the convolution in the same way (dropping the integration limits for notational simplicity):
@@
In the context of convolution, the Dirac delta acts as a form of time translation. When a synaptic kernel $K$ is convolved with a time-shifted delta $\delta(t-a)$, the delta essentially translates the entire shape of the kernel along the timeline to exactly $t=a$. This functionally transforms a complex integral into a simple copy-paste operation, in the sense that for every spike time in the collection of events, the distributional operation pastes a copy of the synaptic response into the total current, allowing discrete events to be encoded into continuous-time dynamics.
@@ 7.4 Distributional differentiation of spike trains
In this section we discuss another property that emphasizes the unique gains of a distribution theory formalism to neuronal spike trains and important topics in neuroscience. Although a spike train has no pointwise derivative, its derivative is well-defined in the distributional sense. By definition, the derivative $s_{i}^{\prime}$ acts on a test function $\varphi$ as $\langle s_{i}^{\prime},\varphi\rangle=-\langle s_{i},\varphi^{\prime}\rangle$, producing $\sum_{k}\varphi^{\prime}(t_{i}^{k})$. In other words, differentiation of a spike train does not measure changes in the spikes themselves, which has no meaning, but rather local changes in the test function at the event times. It tells us how sensitive a downstream process is to the precise timing of spikes.
This is directly relevant to temporal coding hypotheses, where information is carried not by average firing rates but by the precise times at which individual spikes occur. Under a rate code, small shifts in spike timing are noise. But under a temporal code, they are signal. The distributional derivative $\langle s_{i}^{\prime},\varphi\rangle=-\sum_{k}\varphi^{\prime}(t_{i}^{k})$ provides a mathematically exact way to quantify this distinction. It measures how much a given downstream process $\varphi$ changes per unit shift in spike time, without approximation or binning. Temporal coding and its contrast with rate coding have been extensively studied in sensory systems [GerstnerKistlerNaudPaninski2014, theunissen1995temporal, panzeri2010sensory].
This formalism also provides a natural framework for understanding how spike timing jitter affects synaptic integration. In any real neural circuit, spike times are variable, they fluctuate trial to trial due to channel noise, synaptic stochasticity, and network dynamics. The question of whether downstream computation is robust or fragile to this jitter is ultimately a question about the magnitude of $|\varphi^{\prime}(t_{i}^{k})|$ in the sense that if the downstream process changes slowly near the spike time, the computation is robust to small timing perturbations. But if it changes quickly, the computation is necessarily sensitive. This connects the distributional derivative directly to the experimental literature on spike timing precision and its functional consequences [mainen1995reliability, koch1997relationship, galan2008optimal].
Distributional differentiation also provides the formal mathematical basis for analyzing spike-timing-dependent plasticity (STDP), in which the sign and magnitude of synaptic modification depend on the precise temporal order and interval between pre- and postsynaptic spikes, typically within windows of tens of milliseconds. STDP is inherently a derivative-like operation on spike timing. The plasticity rule assigns different outcomes depending on whether a presynaptic spike leads or lags a postsynaptic spike by a small temporal offset. The distributional framework developed here provides the rigorous mathematical structure within which such timing-dependent operations on event-structured objects can be formally defined and analyzed [bi1998synaptic, markram1997regulation, feldman2012spike].
Finally, the distributional derivative supports the analysis of delay-dependent synchronization and refractoriness in neural circuits. In networks where action potentials (or equivalent discrete signals) propagate with finite conduction velocities and neurons experience a refractory period after firing, the causal admissibility of an input depends on the precise timing relationship between signal arrival and the neuron’s refractory state [sila2019, puppo2018]. The derivative of a spike train, acting on a downstream kernel, quantifies exactly how perturbations in these timing relationships propagate through a circuit, a question that cannot be addressed or modeled within a rate-coding or binned-time framework [BrunelHakim1999, ErmentroutTerman2010].
We will develop the intuition of distributional differentiation carefully.
@@
A spike train is a set of instantaneous events, so what does it even mean to differentiate a set like that? When we differentiate ordinary functions, derivatives have physical and geometric interpretations, such as the rate of change, or slope. But distributions like the spike train are not ordinary functions. They have no meaning or values between events, or concepts like slope. This necessitates a change in perspective that ultimately results in a powerful set of mathematical ideas and tools to analyze and understand (in our case) very real world physical and physiological properties of neural signaling and dynamics.
@@
This shift in perspective requires abandoning the question of how to directly differentiate the distribution. Because the spike train has no pointwise values, it is not directly differentiable. Instead, we differentiate the test functions and ask how does the spike train respond.
From the operational requirement formalized in Equation 2.1,
@@
with $s_{i}(t)=\sum_{k}\delta(t-t_{i}^{k})$ (Equation 4.4), the spike train samples the test function at the event times.
@@
This definition is consistent with the ordinary notion of differentiation if $T$ were a smooth function, and preserves integration by parts, but avoids pointwise evaluation. We look at this more deeply (and why the negative sign in the definition needs to be there) below.
@@
The interpretation of Equation 7.11 is that the original spike train $s_{i}$ samples values of the test function $\varphi(t)$ at the event times, while the derivative of the spike train, $s_{i}^{\prime}$, samples values of the derivative of the test function $\varphi^{\prime}(t)$. In other words, $s_{i}^{\prime}$ samples values of the slope of $\varphi(t)$. This means that the derivatives of the spike train (by definition) measures how the test function changes at the event times.
The implication is that $s_{i}^{\prime}$ encodes how sensitive the test function is to timing shifts, i.e., at the spike times. So for example, if $\varphi(t)$ has small amplitudes or is zero near some spike time $t_{i}^{k}$ then $\varphi^{\prime}(t_{i}^{k})\rightarrow 0$ or will be equal to zero at that event time, so the spike (sampling the value $\varphi^{\prime}(t_{i}^{k})$) will produce no response in its derivative. It will not contribute much or anything to the sum in Equation 7.11. If, on the other hand, $\varphi(t)$ changes quickly or steeply near $t_{i}^{k}$, then $|\varphi^{\prime}(t_{i}^{k})|$ will have a large value and the spike at $t_{i}^{k}$ will produce a large contribution to the sum.
Two closing technical comments: First, if $\varphi(t)$ is not differentiable, then $\varphi^{\prime}(t)$ would be undefined, making $s_{i}^{\prime}$ undefined (directly due to its definition). So this operation necessitates all the work we did above to justify why $\varphi(t)\in C_{c}^{\infty}$.
Second, the minus sign in Equation 7.10 results directly from integration by parts. It is worth considering why it arises in some detail, because it provides a deep intuitive insight.
@@
Notice how the right hand side of Equation 7.14d is (by definition) $\langle T^{\prime}_{f},\varphi\rangle$ (c.f. Equation 7.12). The left hand side is Equation 7.13: $\langle T_{f},\varphi^{\prime}\rangle$. Therefore,